# Bayesian Density-Density Regression with Application to Cell-Cell Communications

### v1

@@ Abstract
We introduce a scalable framework for regressing multivariate distributions onto multivariate distributions, motivated by the application of inferring cell-cell communication from population-scale single-cell data. The observed data consist of pairs of multivariate distributions for ligands from one cell type and corresponding receptors from another. For each ordered pair $e=(l,r)$ of cell types $(l\neq r)$ and each sample $i=1,\ldots,n$, we observe a pair of distributions $(F_{ei},G_{ei})$ of gene expressions for ligands and receptors of cell types $l$ and $r$, respectively. The aim is to set up a regression of receptor distributions $G_{ei}$ given ligand distributions $F_{ei}$. A key challenge is that these distributions reside in distinct spaces of differing dimensions. We formulate the regression of multivariate densities on multivariate densities using a generalized Bayes framework with the sliced Wasserstein distance between fitted and observed distributions. Finally, we use inference under such regressions to define a directed graph for cell-cell communications.
@@ 1 Introduction
We propose the first approach for regressing random distributions on random distributions in multivariate settings. A key feature of our approach, which is Bayesian, is its ability to provide a full probabilistic characterization of all relevant unknowns, including data and model parameters. However, this feature can also become a limitation, particularly when dealing with complex relationships in which specifying a complete probabilistic model is challenging. To address this, we adopt a generalized Bayes framework that circumvents the need to fully specify likelihoods. Our approach leverages the sliced Wasserstein distance to measure the discrepancy between the observed response distribution and the predicted distribution, allowing for a principled update of beliefs while maintaining flexibility.
@@
In the usual regression setup, both the predictors and responses are typically scalars or vectors. When distributional variables serve as the predictors and real-valued variables as the response, the density regression problem arises (Póczos et al., , 2013), e.g., predicting health indicators from a distribution of clinical outcomes (Szabó et al., , 2016) or estimating the effects of physical activity on biomarkers (Matabuena and Petersen, , 2023). Another type of density regression arises with real-valued predictors and density-valued responses. This is studied, for example, in Tokdar et al., (2004); Dunson et al., (2007) or Shen and Ghosal, (2016). Recently, the density-density regression (DDR) problem, which deals with distributional predictor and distributional response variables, has been studied. For example, regression with one-dimensional predictor and one-dimensional response is considered in Zhao et al., (2023) using a Riemannian representation of density functions. The method of Zhao et al., (2023) is hard to extend to multivariate settings due to the usage of a wrapping function, which is only well-defined in one dimension. To the best of our knowledge, there are currently no methods for DDR with multivariate distributions.
@@
We propose a DDR model, where both predictors and responses are random distributions of arbitrary dimensions. The motivation comes from the challenge of modeling molecular dependence relationships among different cell types in population-scale single-cell data. In this type of data, gene expressions are measured for a large number of cells donated by a large number of subjects. Exploiting that cell types can be identified based on known cellular markers, this provides an opportunity to decipher the communication patterns between different cell types by way of expressions of ligand-receptor pairs. Because of the availability of a large number of cells from each donor, we, in essence, observe for each subject the empirical distributions of the gene expressions for each cell type. Therefore, we can cast the inference problem of cell-cell communications as a DDR problem, where the predictor for each subject is the distribution of gene expression for the ligands from one cell type and the response is the distribution for receptors from another cell type. Conventional bulk gene expression methods collapse the distributional representation by taking the mean as the representative statistic for downstream regression. In contrast, the proposed Bayesian DDR approach allows us to model complex regression dependencies between distributions while preserving the full variability and uncertainty of the data.
@@
Our inference procedure is based on the posterior distribution for a regression function that maps from the space of the predictor distribution to the space of the response distribution. Avoiding an explicit sampling model, we rely instead on a generalized likelihood, which is constructed based on a discrepancy between the pushforward (predicted) response distribution and the observed response distribution. Hence, the proposed Bayesian DDR approach falls within the generalized Bayesian inference framework (Bissiri et al., , 2016).
@@
One challenge in implementing DDR is that we do not directly observe distributions through their density functions. Instead, we observe finite samples from distributions. Moreover, the number of observed samples can differ between distributions, and distributions may lie in high-dimensional spaces. Therefore, for the construction of the generalized likelihood we need to use a discrepancy between distributions that can be reliably estimated with finite samples in high dimensions. We choose the sliced Wasserstein distance (Rabin et al., , 2012; Bonneel et al., , 2015; Peyré et al., , 2019), which is a sliced variant (random projection variant) of the Wasserstein distance (Villani, , 2009). Sliced Wasserstein does not suffer from the curse of dimensionality and is computationally scalable with respect to the number of support points in the distributions. Moreover, the use of sliced Wasserstein allows us to derive a Metropolis-adjusted Langevin algorithm (MALA, Girolami and Calderhead, 2011) for efficient inference.
@@
Using the proposed Bayesian DDR, we construct a cell-cell communication network of different cell types based on ligand–receptor interactions for each pair of cell types. We perform Bayesian DDR for each pair of cell types, i.e., regressing distributions over receptors onto distributions over ligands. After fitting the Bayesian DDR models, we propose two ways to infer a network. The first option is a weighted fully-connected graph with edge weights derived from the expected relative errors of the regressions. The second approach casts the graph selection as a decision problem, which we implement by way of a simple decision boundary. In particular, we parameterize the regression as a linear function and put a shrinkage prior on the linear coefficient. We report edges by thresholding the posterior probabilities in these linear functions, with the threshold determined by minimizing posterior expected false negative rate (FNR) under a bound on posterior expected false discovery rate (FDR).
@@
The remainder of the article is organized as follows. In Section 2, we review the definitions of Wasserstein and sliced Wasserstein distance, and their statistical and computational properties. In Section 3, we propose a generalized Bayesian DDR framework using sliced Wasserstein distance to define a generalized likelihood, and we derive posterior simulation using a MALA sampler. In Section 4, we apply the proposed Bayesian DDR approach to infer cell-cell communication between two cell types. In Section 5, we build on the regression setup and propose discovery of a directed graph by way of forming Bayesian DDRs for ordered pairs of cell types. In Section 6, we apply the proposed graph discovery for graphs among multiple cell types in a population-scale single-cell dataset, and we validate inference in a semi-simulation with real predictor distributions and a simulated true regression function.
For notations, we use $\delta_{x}$ to denote the Dirac delta function, which takes the value 1 at $x$ and 0 otherwise. For any $d\geq 2$, we denote $\mathbb{S}^{d-1}=\{\theta\in\mathbb{R}^{d}\mid||\theta||_{2}^{2}=1\}$ as the unit hypersphere. We use $I$ for the identity matrix and $\mathbf{1}$ for the vector of all 1’s, with the size depending on the context. For any two sequences $a_{n}$ and $b_{n}$, the notation $a_{n}=\mathcal{O}(b_{n})$ means that $a_{n}\leq Cb_{n}$ for all $n\geq 1$, where $C$ is some universal constant. Other notations will be introduced later when they are used.
@@ 2 Background on Sliced Wasserstein Distance
By way of a brief review of sliced Wasserstein distance, we introduce some notation and definitions. Given $p\geq 1$, let $G_{1},G_{2}\in\mathcal{P}_{p}(\mathbb{R}^{d})$ where $\mathcal{P}_{p}(\mathbb{R}^{d})$ be the set of all distributions supported on the $\mathbb{R}^{d}$ with finite $p$-th moment. Wasserstein-$p$ ($p\geq 1$) distance (Villani, , 2009) between $G_{1}$ and $G_{2}$ is defined as:
@@
where $\Pi(G_{1},G_{2})=\left\{\pi\in\mathcal{P}(\mathbb{R}^{d}\times\mathbb{R}^{d})\mid\pi(A,\mathbb{R}^{d})=G_{1}(A),\ \pi(\mathbb{R}^{d},B)=G_{2}(B)\ \forall A,B\subset\mathbb{R}^{d}\right\}$ is the set of all transportation plans/couplings. When observing $x_{1},\ldots,x_{m_{1}}\overset{i.i.d.}{\sim}G_{1}$ and $y_{1},\ldots,y_{m_{2}}\overset{i.i.d.}{\sim}G_{2}$, the plug-in estimation of the Wasserstein-$p$ distance is given by:
$W_{p}^{p}({G}_{1,m_{1}},{G}_{2,m_{2}})=\min_{\gamma\in\Gamma(\mathbf{1}/m_{1},\mathbf{1}/m_{2})}\sum_{i=1}^{m_{1}}\sum_{j=1}^{m_{2}}\|x_{i}-y_{j}\|_{p}^{p}\gamma_{ij},$ (2)
@@
where ${G}_{1,m_{1}}=\frac{1}{m_{1}}\sum_{i=1}^{m_{1}}\delta_{x_{i}}$ and ${G}_{2,m_{2}}=\frac{1}{m_{2}}\sum_{j=1}^{m_{2}}\delta_{y_{j}}$ are corresponding empirical distributions over i.i.d. samples from $G_{1}$ and $G_{2}$, and the set of discrete transportation plans is $\Gamma(\mathbf{1}/m_{1},\mathbf{1}/m_{2})=\left\{\gamma\in\mathbb{R}_{+}^{m_{1}\times m_{2}}\mid\gamma\mathbf{1}=\mathbf{1}/m_{1}\,\gamma^{\top}\mathbf{1}=\mathbf{1}/m_{2}\right\}$. Computing $W_{p}^{p}({G}_{1,m_{1}},{G}_{2,m_{2}})$ is a linear programming problem, which costs $\mathcal{O}((m_{1}+m_{2})^{3}\log(m_{1}+m_{2}))$ in time complexity (Peyré et al., , 2019). From Fournier and Guillin, (2015), we have the following sample complexity: $\mathbb{E}\left[\left|W_{p}({G}_{1,m_{1}},{G}_{2,m_{2}})-W_{p}(G_{1},G_{2})\right|\right]=\mathcal{O}(m_{1}^{-1/d}+m_{2}^{-1/d})$, which implies that we need $m_{1}$ and $m_{2}$ to be large to achieve small approximation errors. However, the time complexity to compute the Wasserstein distance scales poorly with $m_{1}$ and $m_{2}$.
@@
This conundrum is addressed by introducing sliced Wasserstein (SW) distance (Rabin et al., , 2014; Bonneel et al., , 2015). SW relies on the fact that one-dimensional Wasserstein distance can be computed efficiently. In particular, let $G_{1},G_{2}\in\mathcal{P}_{p}(\mathbb{R})$, computing $W_{p}^{p}({G}_{1,m_{1}},{G}_{2,m_{2}})$ in Equation (2) costs only $\mathcal{O}((m_{1}+m_{2})\log(m_{1}+m_{2}))$ in time complexity by using the north-west corner solution (Peyré et al., , 2019).
@@
SW distance between two distributions $G_{1},G_{2}\in\mathcal{P}_{p}(\mathbb{R}^{d})$ is defined as: $SW_{p}^{p}(G_{1},G_{2})=\mathbb{E}_{\theta\sim\mathcal{U}(\mathbb{S}^{d-1})}[W_{p}^{p}(\theta\sharp G_{1},\theta\sharp G_{2})],$ where $\mathcal{U}(\mathbb{S}^{d-1})$ is the uniform distribution over the unit hypersphere in $d$ dimension, and $\theta\sharp G_{1}$ and $\theta\sharp G_{2}$ denote the pushforward distribution of $G_{1}$ and $G_{2}$ through the function $f_{\theta}(x)=\theta^{\top}x$. In particular, given two measurable spaces $(\mathcal{X}_{1},\Sigma_{1})$ and $(\mathcal{X}_{2},\Sigma_{2})$, a measurable function $f:\mathcal{X}_{1}\to\mathcal{X}_{2}$, and a measure $\mu:\Sigma_{1}\to[0,\infty)$, the pushforward of $\mu$ through $f$ is $f\sharp\mu(B)=\mu(f^{-1}(B))$ for any $B\in\Sigma_{2}$ – we will use the definition with probability measures only. Similar to Wasserstein, we have the following plug-in estimation of SW:
$SW_{p}^{p}({G}_{1,m_{1}},{G}_{2,m_{2}})=\mathbb{E}_{\theta\sim\mathcal{U}(\mathbb{S}^{d-1})}[W_{p}^{p}(\theta\sharp{G}_{1,m_{1}},\theta\sharp{G}_{2,m_{2}})].$ (3)
@@
From Nadjahi et al., (2020); Nguyen et al., (2021); Nietert et al., (2022); Boedihardjo, (2025), we have that $\mathbb{E}\left[\left|SW_{p}({G}_{1,m_{1}},{G}_{2,m_{2}})-SW_{p}(G_{1},G_{2})\right|\right]=\mathcal{O}(m_{1}^{-1/2}+m_{2}^{-1/2})$. In particular, there is no dimension dependence. Combining with the linear-time complexity of empirical SW, we have a both, computationally and statistically, scalable solution for comparing distributions. The last step to compute SW is approximating the expectation with respect to $\mathcal{U}(\mathbb{S}^{d-1})$ (Nguyen et al., , 2024; Leluc et al., , 2024; Nguyen and Ho, , 2024; Sisouk et al., , 2025). For example, we can use a simple Monte Carlo estimation:
$\widehat{SW}_{p}^{p}({G}_{1,m_{1}},{G}_{2,m_{2}};L)=\frac{1}{L}\sum_{l=1}^{L}W_{p}^{p}(\theta_{l}\sharp{G}_{1,m_{1}},\theta_{l}\sharp{G}_{2,m_{2}}),$ (4)
@@
where $\theta_{1},\ldots,\theta_{L}\overset{i.i.d.}{\sim}\mathcal{U}(\mathbb{S}^{d-1})$ with $L$ being the number of Monte Carlo samples or the number of projections. The overall time complexity for computing this approximation is $\mathcal{O}(L(m_{1}+m_{2})\log(m_{1}+m_{2})+Ld(m_{1}+m_{2}))$ where $\mathcal{O}(Ld(m_{1}+m_{2}))$ is for projecting supports of ${G}_{1,m_{1}}$ and ${G}_{2,m_{2}}$, and $\mathcal{O}(L(m_{1}+m_{2})\log(m_{1}+m_{2}))$ is for computing $W_{p}^{p}(\theta_{l}\sharp{G}_{1,m_{1}},\theta_{l}\sharp{G}_{2,m_{2}})$ with $\theta_{l}\sharp{G}_{1,m_{1}}=\frac{1}{m_{1}}\sum_{i=1}^{m_{1}}\delta_{\theta_{l}^{\top}x_{i}}$ and $\theta_{l}\sharp{G}_{2,m_{2}}=\frac{1}{m_{2}}\sum_{j=1}^{m_{2}}\delta_{\theta_{l}^{\top}y_{j}}$ for $l=1,\ldots,L$
3 Generalized Bayesian Density-Density Regression
3.1 Generalized Bayesian Density-Density Regression
We consider the inference problem of relating a distribution-valued response $G_{i}$ to a distribution-valued predictor $F_{i}$. We observe samples $x_{i1},\ldots,x_{im_{i1}}\sim F_{i}\in\mathcal{P}_{2}(\mathbb{R}^{d_{1}})$ and $y_{i1},\ldots,y_{im_{i2}}\sim G_{i}\in\mathcal{P}_{2}(\mathbb{R}^{d_{2}})$ for $i=1,\ldots,N$. Let $\mathcal{S}=\{(\{x_{i1},\ldots,x_{im_{i1}}\},\{y_{i1},\ldots,y_{im_{i2}}\})\}_{i=1}^{N}$ denote the observed data where $x_{i1},\ldots,x_{im_{i1}}\in\mathbb{R}^{d_{1}}$ and $y_{i1},\ldots,y_{im_{i2}}\in\mathbb{R}^{d_{2}}$. Equivalently, we represent the data as $\mathcal{S}=\{(\widehat{F}_{i},\widehat{G}_{i})\}_{i=1}^{N}$ where $\widehat{F}_{i}=\frac{1}{m_{i1}}\sum_{j=1}^{m_{i1}}\delta_{x_{ij}}$ and $\widehat{G}_{i}=\frac{1}{m_{i2}}\sum_{j=1}^{m_{i2}}\delta_{y_{ij}}$ are the empirical distributions. Assuming large enough sample sizes $m_{i1}$ and $m_{i2}$, we drop the distinction of $({F}_{i},{G}_{i})$ versus $(\widehat{F}_{i},\widehat{G}_{i})$.
We define the following DDR model:
${G}_{i}|{F}_{i},f\sim p({G}_{i}\mid{F}_{i},f)\propto\exp\left(-w\sum_{i=1}^{N}\widehat{SW}_{2}^{2}(f\sharp{F}_{i},{G}_{i};L)\right),$ (5)

### v2

# Bayesian Multivariate Density-Density Regression

@@ Abstract
We introduce a novel and scalable Bayesian framework for multivariate-density-density regression (DDR), designed to model relationships between multivariate distributions. Our approach addresses the critical issue of distributions residing in spaces of differing dimensions. We utilize a generalized Bayes framework, circumventing the need for a fully specified likelihood by employing the sliced Wasserstein distance to measure the discrepancy between fitted and observed distributions. This choice not only handles high-dimensional data and varying sample sizes efficiently but also facilitates a Metropolis-adjusted Langevin algorithm (MALA) for posterior inference. Furthermore, we establish the posterior consistency of our generalized Bayesian approach, ensuring that the posterior distribution concentrates around the true parameters as the sample size increases. Through simulations and application to a population-scale single-cell dataset, we show that Bayesian DDR provides robust fits, superior predictive performance compared to traditional methods, and valuable insights into complex biological interactions.
@@ 1 Introduction
We propose the first approach for regressing random distributions on random distributions in multivariate settings. A key feature of our approach, which is Bayesian, is its ability to provide a full probabilistic characterization of all relevant unknowns, including data and model parameters. However, this feature can also become a limitation, particularly when dealing with complex relationships in which specifying a complete probabilistic model is challenging. To address this, we adopt a generalized Bayes framework that circumvents the need to start with a fully specify likelihood. Our approach leverages the sliced Wasserstein distance to measure the discrepancy between the observed response distribution and a predicted distribution, allowing for a principled update of beliefs while maintaining flexibility.
@@
In the usual regression setup, both the predictors and responses are typically scalars or vectors. When distributional variables serve as the predictors and real-valued variables as the response, the density regression problem arises (Póczos et al.,, 2013), e.g., predicting health indicators from a distribution of clinical outcomes (Szabó et al.,, 2016) or estimating the effects of physical activity on biomarkers (Matabuena and Petersen,, 2023). Another type of density regression arises with real-valued predictors and density-valued responses. This is studied, for example, in Tokdar et al., (2004); Dunson et al., (2007) or Shen and Ghosal, (2016). Recently, the density-density regression (DDR) problem, which deals with distributional predictor and distributional response variables, has been studied. For example, regression with one-dimensional predictor and one-dimensional response is considered in Zhao et al., (2023) using a Riemannian representation of density functions. The method of Zhao et al., (2023) is hard to extend to multivariate density functions due to the usage of a wrapping function, which is only well-defined in one dimension. To the best of our knowledge, there are currently no methods for DDR with multivariate distributions. Extending existing univariate methods to this multivariate setting is non-trivial, as most of them depend on the notion of the quantile function, which lacks a universally accepted definition for multivariate distributions.
@@
We propose a DDR model, where both predictors and responses are random distributions of arbitrary dimensions. The motivation comes from the challenge of modeling molecular dependence relationships among different cell types in population-scale single-cell data. In this type of data, gene expressions are measured for a large number of cells donated by a large number of subjects. Exploiting that cell types can be identified based on known cellular markers, this provides an opportunity to decipher the communication patterns between different cell types by way of expressions of ligand-receptor pairs. Because of the availability of a large number of cells from each donor, we, in essence, observe for each subject the empirical distributions of the gene expressions for each cell type. Therefore, we can cast the inference problem of cell-cell communications as a DDR problem, where the predictor for each subject is the distribution of gene expression for the ligands from one cell type (sender cells produce ligands as signaling molecules), and the response is the distribution of receptors from another cell type (receiver cells express receptors and respond to the signals). Conventional bulk gene expression methods collapse the distributional representation by taking the mean as the representative statistic for downstream regression. In contrast, the proposed Bayesian DDR approach allows us to model complex regression dependencies between distributions while preserving the full variability and uncertainty of the data.
@@
Our inference procedure is based on the posterior distribution for a regression function that maps from the space of the predictor distribution to the space of the response distribution. Avoiding to state an explicit sampling model, we rely instead on a generalized likelihood, which is constructed based on a discrepancy between the pushforward (predicted) response distribution and the observed response distribution. Hence, the proposed Bayesian DDR approach falls within the generalized Bayesian inference framework (Bissiri et al.,, 2016).
@@
One challenge in implementing DDR is that we do not directly observe distributions through their density functions. Instead, we observe finite samples from distributions. Moreover, the number of observed samples can differ between distributions, and distributions may lie in high-dimensional spaces. Therefore, we base the construction of the generalized likelihood on a discrepancy between distributions that can be reliably estimated with finite samples in high dimensions. We choose the sliced Wasserstein distance (Rabin et al.,, 2012; Bonneel et al.,, 2015; Peyré et al.,, 2019), which is a sliced variant of the Wasserstein distance through random projection  (Villani,, 2009). Sliced Wasserstein does not suffer from the curse of dimensionality and is computationally scalable with respect to the number of support points in the distributions. Moreover, the use of sliced Wasserstein allows us to derive a Metropolis-adjusted Langevin algorithm (MALA, Girolami and Calderhead, 2011) for efficient posterior inference and to guarantee the consistency of the posterior.
@@
Using the proposed Bayesian DDR, we construct a directed cell-cell communication network of different cell types based on ligand–receptor interactions for each ordered pair of cell types. We carry out Bayesian DDR for each pair of cell types, i.e., regressing distributions over receptors onto distributions over ligands, implemented as a separate instance of DDR for each pair of cell types. After fitting the Bayesian DDR models, we propose two ways to infer the cell-cell communication network. The first option is a weighted fully-connected graph with edge weights derived from the expected relative errors of the regressions. The second approach casts the graph selection as a decision problem, which we implement by way of a simple decision boundary. For simplicity, we parameterize the regression as a linear function and put a shrinkage prior on the linear coefficient, which can be easily extended to nonlinear functions. We report edges by thresholding the posterior probabilities in these linear functions, with the threshold determined by minimizing posterior expected false negative rate (FNR) under a bound on posterior expected false discovery rate (FDR).
@@
The remainder of the article is organized as follows. In Section 2, we review the definitions of Wasserstein and sliced Wasserstein distance, and their statistical and computational properties. In Section 3, we propose a generalized Bayesian DDR framework using sliced Wasserstein distance to define a generalized likelihood, and we derive posterior simulation using a MALA sampler. In Section 4, we apply the proposed Bayesian DDR approach to infer cell-cell communication between two cell types. In Section 5, we build on the regression setup and propose discovery of a directed graph by way of forming Bayesian DDRs for ordered pairs of cell types. In Section 6, we apply the proposed graph discovery for graphs among multiple cell types in a population-scale single-cell dataset, and we validate inference in a semi-simulation with real predictors and a simulated true regression function.
[…]
@@ 2 Background on Sliced Wasserstein Distance
By way of a brief review of sliced Wasserstein distance, we introduce some notation and definitions. Given $p\geq 1$, let $G_{1},G_{2}\in\mathcal{P}_{p}(\mathbb{R}^{d})$ where $\mathcal{P}_{p}(\mathbb{R}^{d})$ be the set of all distributions supported on $\mathbb{R}^{d}$ with finite $p$-th moment. Wasserstein-$p$ ($p\geq 1$) distance (Villani,, 2009) between $G_{1}$ and $G_{2}$ is defined as:
@@
where $\Pi(G_{1},G_{2})=\left\{\pi\in\mathcal{P}(\mathbb{R}^{d}\times\mathbb{R}^{d})\mid\pi(A,\mathbb{R}^{d})=G_{1}(A),\ \pi(\mathbb{R}^{d},B)=G_{2}(B)\ \forall A,B\subset\mathbb{R}^{d}\right\}$ is the set of all transportation plans/couplings. When observing $x_{1},\ldots,x_{m_{1}}\overset{i.i.d.}{\sim}G_{1}$ and $y_{1},\ldots,y_{m_{2}}\overset{i.i.d.}{\sim}G_{2}$, a plug-in estimation of the Wasserstein-$p$ distance is given by:
$W_{p}^{p}(\widehat{G}_{1},\widehat{G}_{2})=\min_{\gamma\in\Gamma(\mathbf{1}/m_{1},\mathbf{1}/m_{2})}\sum_{i=1}^{m_{1}}\sum_{j=1}^{m_{2}}\|x_{i}-y_{j}\|_{p}^{p}\gamma_{ij},$ (2)
@@
where $\widehat{G}_{1}=\frac{1}{m_{1}}\sum_{i=1}^{m_{1}}\delta_{x_{i}}$ and $\widehat{G}_{2}=\frac{1}{m_{2}}\sum_{j=1}^{m_{2}}\delta_{y_{j}}$ are corresponding empirical distributions over i.i.d. samples from $G_{1}$ and $G_{2}$, and the set of discrete transportation plans is $\Gamma(\mathbf{1}/m_{1},\mathbf{1}/m_{2})=\left\{\gamma\in\mathbb{R}_{+}^{m_{1}\times m_{2}}\mid\gamma\mathbf{1}=\mathbf{1}/m_{1}\text{ and }\gamma^{\top}\mathbf{1}=\mathbf{1}/m_{2}\right\}$. Computing $W_{p}^{p}(\widehat{G}_{1},\widehat{G}_{2})$ is a linear programming problem, which costs $\mathcal{O}((m_{1}+m_{2})^{3}\log(m_{1}+m_{2}))$ in time complexity (Peyré et al.,, 2019). From Fournier and Guillin, (2015), we have the following sample complexity: $\mathbb{E}\left[\left|W_{p}(\widehat{G}_{1},\widehat{G}_{2})-W_{p}(G_{1},G_{2})\right|\right]=\mathcal{O}(m_{1}^{-1/d}+m_{2}^{-1/d})$, which implies that we need $m_{1}$ and $m_{2}$ to be large to achieve small approximation errors. However, the time complexity to compute the Wasserstein distance scales poorly with $m_{1}$ and $m_{2}$.
@@
This conundrum is addressed by introducing sliced Wasserstein (SW) distance (Rabin et al.,, 2014; Bonneel et al.,, 2015). SW relies on the fact that one-dimensional Wasserstein distance can be computed efficiently. In particular, let $G_{1},G_{2}\in\mathcal{P}_{p}(\mathbb{R})$, computing $W_{p}^{p}(\widehat{G}_{1},\widehat{G}_{2})$ in Equation (2) costs only $\mathcal{O}((m_{1}+m_{2})\log(m_{1}+m_{2}))$ in time complexity by using the north-west corner solution (Peyré et al.,, 2019).
SW distance between two distributions $G_{1},G_{2}\in\mathcal{P}_{p}(\mathbb{R}^{d})$ is defined as:
$SW_{p}^{p}(G_{1},G_{2})=\mathbb{E}_{\theta\sim\mathcal{U}(\mathbb{S}^{d-1})}[W_{p}^{p}(\theta\sharp G_{1},\theta\sharp G_{2})],$ (3)
@@
where $\mathcal{U}(\mathbb{S}^{d-1})$ is the uniform distribution over the unit hypersphere in $d$ dimension, and $\theta\sharp G_{1}$ and $\theta\sharp G_{2}$ denote the pushforward distribution of $G_{1}$ and $G_{2}$ through the function $f_{\theta}(x)=\theta^{\top}x$. In particular, given two measurable spaces $(\mathcal{X}_{1},\Sigma_{1})$ and $(\mathcal{X}_{2},\Sigma_{2})$, a measurable function $f:\mathcal{X}_{1}\to\mathcal{X}_{2}$, and a measure $\mu:\Sigma_{1}\to[0,\infty)$, the push-forward of $\mu$ through $f$ is $f\sharp\mu(B)=\mu(f^{-1}(B))$ for any $B\in\Sigma_{2}$ – we will use the definition with probability measures only. The push-forward measure is a general measure-theoretic concept. When the mapping is bijective, typically between spaces of the same dimension, the change of variables formula can be used to explicitly compute the density of the push-forward measure, provided the measures involved are absolutely continuous. Similar to Wasserstein, we have the following plug-in estimation of SW:
$SW_{p}^{p}(\widehat{G}_{1},\widehat{G}_{2})=\mathbb{E}_{\theta\sim\mathcal{U}(\mathbb{S}^{d-1})}[W_{p}^{p}(\theta\sharp\widehat{G}_{1},\theta\sharp\widehat{G}_{2})].$ (4)
@@
From Nadjahi et al., (2020); Nguyen et al., (2021); Nietert et al., (2022); Boedihardjo, (2025), we have that $\mathbb{E}\left[\left|SW_{p}(\widehat{G}_{1},\widehat{G}_{2})-SW_{p}(G_{1},G_{2})\right|\right]=\mathcal{O}(m_{1}^{-1/2}+m_{2}^{-1/2})$. In particular, there is no dimension dependence. Combining with the linear-time complexity of empirical SW, we have a both, computationally and statistically, scalable solution for comparing distributions. The last step to compute SW is approximating the expectation with respect to $\mathcal{U}(\mathbb{S}^{d-1})$ (Nguyen et al.,, 2024; Leluc et al.,, 2024; Nguyen and Ho,, 2024; Sisouk et al.,, 2025). For example, we can use simple Monte Carlo estimation:
$\widehat{SW}_{p}^{p}(\widehat{G}_{1},\widehat{G}_{2};L)=\frac{1}{L}\sum_{l=1}^{L}W_{p}^{p}(\theta_{l}\sharp\widehat{G}_{1},\theta_{l}\sharp\widehat{G}_{2}),$ (5)
@@
where $\theta_{1},\ldots,\theta_{L}\overset{i.i.d.}{\sim}\mathcal{U}(\mathbb{S}^{d-1})$ with $L$ being the number of Monte Carlo samples or the number of projections. The overall time complexity for computing this approximation is $\mathcal{O}(L(m_{1}+m_{2})\log(m_{1}+m_{2})+Ld(m_{1}+m_{2}))$ where $\mathcal{O}(Ld(m_{1}+m_{2}))$ is for projecting supports of $\widehat{G}_{1}$ and $\widehat{G}_{2}$, and $\mathcal{O}(L(m_{1}+m_{2})\log(m_{1}+m_{2}))$ is for computing $W_{p}^{p}(\theta_{l}\sharp\widehat{G}_{1},\theta_{l}\sharp\widehat{G}_{2})$ with $\theta_{l}\sharp\widehat{G}_{1}=\frac{1}{m_{1}}\sum_{i=1}^{m_{1}}\delta_{\theta_{l}^{\top}x_{i}}$ and $\theta_{l}\sharp\widehat{G}_{2}=\frac{1}{m_{2}}\sum_{j=1}^{m_{2}}\delta_{\theta_{l}^{\top}y_{j}}$ for $l=1,\ldots,L$ .
3 Generalized Multivariate Bayesian Density-Density Regression
3.1 Generalized Multivariate Bayesian Density-Density Regression
We consider the inference problem of relating a distribution-valued response $G_{i}\in\mathcal{P}_{2}(\mathbb{R}^{d_{2}})$ ($d_{2}\geq 2$) to a distribution-valued predictor $F_{i}\in\mathcal{P}_{2}(\mathbb{R}^{d_{1}})$ ($d_{1}\geq 2$) for $i=1,\ldots,N$ ($N>0$). We represent the data as $\mathcal{S}=\{(F_{i},G_{i})\}_{i=1}^{N}$ and define the following DDR model. The model is specified as a generalized likelihood (Bissiri et al.,, 2016) $\ell(f;{G}_{i},{F}_{i})$ based on a loss function for a fitted approximation of ${G}_{i}$:
$\ell(f;{G}_{i},{F}_{i})=\exp\left(-wSW_{2}^{2}(f\sharp{F}_{i},{G}_{i})\right),$ (6)